# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 7.2. Rank two examples
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.7, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by blowing up three smooth points on the three components of the toric boundary divisor $\overline{D}$ of $\overline{Y}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.7 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of a toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$. It follows from the dual web represented on the right of Figure 7.7 that the central fiber $\mathcal{X}_{0}$ contains two compact irreducible components. The compact component corresponding to the triangle in the web is isomorphic to $\mathbb{P}^{2}$. On the other hand, the compact component corresponding to the other bounded cell of the web is isomorphic to the Hirzebruch surface $\mathbb{F}_{6}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$. The associated rank two 5d SCFT appears as $\mathbb{P}^{2}\cup\mathbb{F}_{6}$ in [63, Figure 7].
[…]
Figure 7.7. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering a rank two 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 7.8, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by blowing up four smooth points on the four components of the toric boundary divisor $\overline{D}$ of $\overline{Y}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 7.8 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of a toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$. It follows from the dual web represented on the right of Figure 7.8 that the central fiber $\mathcal{X}_{0}$ contains two compact irreducible components. The compact component corresponding to the quadrilatera in the web is isomorphic to $\mathbb{F}_{1}$. On the other hand, the compact component corresponding to the other bounded cell of the web is isomorphic to $\mathbb{F}_{6}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$. The associated rank two 5d SCFT appears as $\mathbb{F}_{1}\cup\mathbb{F}_{6}$ in [63, Figure 8], where it is also identified with the 5d $\mathcal{N}=1$ $\mathrm{Sp}(2)_{0}$ gauge theory.
[…]
Figure 7.8. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering a rank two 5d SCFT.
@@ 7.3. Relation to the Fanosearch program
It is shown in [30, Theorem 39] that there exists actually a one-to-one correspondence between the set of $\mathbb{Q}$-Gorenstein degenerations of the Fano surface $X_{\overline{P}}$ to Fano toric surfaces and the set of Fano polygons obtained from $\overline{P}$ by mutations. Finally, one can attach to a Fano polygon $\overline{P}$ a log Calabi–Yau surface $(Y,D)$. For this, denote by $(\overline{Y},\overline{D},\overline{L})$ the polarized toric variety with momentum polytope $\overline{P}$. Then, $(Y,D)$ is obtained from $(\overline{Y},\overline{D})$ by blowing up for each elementary T-cone of $\overline{P}$ a point on the corresponding toric divisor of $\overline{Y}$. Finally, one can define a line bundle $L$ on $(Y,D)$ by $L=p^{\star}\overline{L}\otimes\mathcal{O}(-\sum_{i}a_{i}E_{i})$, where $p:Y\rightarrow\overline{Y}$ is the toric morphism, $E_{i}$ are the exceptional curves, and $a_{i}$ are the common lattice height/base of the corresponding elementary T-cones. The set of cluster charts $(\mathbb{C}^{\star})^{2}\subset U=Y\setminus D$ is in one-to-one correspondence with the set of mutations of $\overline{P}$, and the space of sections $H^{0}(Y,L)$ can be identified with the set of maximally mutable Laurent polynomials defined in [3, Definition 4], which are expected to be Landau-Ginzburg mirror to the Fano surface $X_{\overline{P}}$.
@@
The set-up of mirror symmetry for Fano surfaces briefly reviewed above can be viewed as a special case of the set-up considered in the present paper. Indeed, if $(Y,D,L)$ is a log Calabi–Yau surface with line bundle obtained from a Fano polygon $\overline{P}$, then, we have $L=\mathcal{O}(\sum_{i}a_{i}D_{i})$, where $D_{i}$ are the irreducible components of $D$, and $a_{i}$ is the lattice distance between the origin 00 and the side of $\overline{P}$ corresponding to $D_{i}$. In other words, we are in the particular case where the divisors defined by $L$ are linearly equivalent to a divisor supported on $D$. Conversely, if $(Y,D,L)$ is a log Calabi–Yau surface such that $L=\mathcal{O}(\sum_{i}a_{i}D_{i})$ with $a_{i}\in\mathbb{Z}_{\geq 0}$, then, for every toric model given by a polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ with momentum polytope $\overline{P}$, there exists a torus-invariant section of $\overline{L}$ with zero-divisor $\sum_{i}a_{i}\overline{D}_{i}$, which corresponds to a particular integral point in $\overline{P}$ which can be taken as the origin [74]. In this particular situation, the Fano polygon $\overline{P}$ naturally defines a Symington polygon $P$ by cutting out all elementary T-cones, so that the origin is the only singularity of the integral affine structure, see Figure 7.9. The corresponding non-compact Calabi–Yau 3-fold is the total space of the canonical line bundle $K_{X_{\overline{P}}}$ of the Fano surface $X_{\overline{P}}$. Indeed, starting from the toric Calabi–Yau 3-fold $K_{\overline{X}_{\overline{P}}}$, and applying the non-toric deformation described in §6.2, the Fano toric surface $\overline{X}_{\overline{P}}$ is deformed non-torically into the Fano surface $X_{\overline{P}}$. In particular, the corresponding canonical 3-fold singularity $\mathcal{X}^{\mathrm{can}}$ is obtained from $K_{X_{\overline{P}}}$ by contracting the zero section to a point. The 3-fold $K_{X_{\overline{P}}}$ is singular in general and so is only a partial crepant resolution of $\mathcal{X}^{\mathrm{can}}$. To obtain a crepant resolution, one should resolve the singularities by decomposing the Symington polygon $P$ into triangles of size one.
[…]
@@
Figure 7.9. On the left, a Fano polygon with a decomposition into elementary T-cones and R-cones. On the right, the corresponding Symington polygon obtained by cutting out the T-cones.
@@
As shown in [57], all $\mathbb{Q}$-Gorenstein degenerations of $\mathbb{P}^{2}$ are weighted projective spaces $\mathbb{P}(a^{2},b^{2},c^{2})$, for Markov triples $(a,b,c)$, that is, triple of positive integers such that $a^{2}+b^{2}+c^{2}=3abc$. Correspondingly, all mutation equivalent Fano polygons to the Fano polygon $\overline{P}$ in Example 7.1, can be characterized by a mutation sequence indexed by Markov triples as explained in [65, Example 2.4]. We illustrate in Figure 7.10 one of these mutation equivalent polygons $\overline{P}^{\prime}$, along with a polyhedral decomposition which determines a degeneration of the toric variety $(\overline{Y}^{\prime},\overline{D}^{\prime})$ with momentum polytope image $\overline{P}^{\prime}$. The mirror $\mathcal{X}\to\mathbb{C}$ to this degeneration is a deformation of the toric Calabi–Yau 3-fold $K_{\mathbb{P}(1,1,4)}$, and is still isomorphic to $K_{\mathbb{P}^{2}}$, as in Example 7.1. The compact component $\mathbb{P}^{2}$ of $\mathcal{X}_{0}$ is obtained from the compact component $\mathbb{P}(1,1,4)$ of $\overline{\mathcal{X}}_{0}$ by locally smoothing as in Equation (6.2) the elementary T-singularity $\mathbb{C}^{2}/(\mathbb{Z}/4\mathbb{Z})$ corresponding to the elementary T-cone drawn in red in Figure 7.10.
[…]
Figure 7.10. The momentum polytope $\overline{P}^{\prime}$ for a polarized toric variety $\overline{Y}^{\prime}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}^{\prime}$ of $\overline{P}^{\prime}$ defining a maximal degeneration of $\overline{Y}^{\prime}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{0}$ 5d SCFT.

### v2

@@ 7.2. Rank two examples
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 34, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by blowing up three smooth points on the three components of the toric boundary divisor $\overline{D}$ of $\overline{Y}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 34 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of a toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$. It follows from the dual web represented on the right of Figure 34 that the central fiber $\mathcal{X}_{0}$ contains two compact irreducible components. The compact component corresponding to the triangle in the web is isomorphic to $\mathbb{P}^{2}$. On the other hand, the compact component corresponding to the other bounded cell of the web is isomorphic to the Hirzebruch surface $\mathbb{F}_{6}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$. The associated rank two 5d SCFT appears as $\mathbb{P}^{2}\cup\mathbb{F}_{6}$ in [66, Figure 7].
Figure 34. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering a rank two 5d SCFT.
@@
Let $\overline{P}$ be the decorated lattice polygon shown on the left of Figure 35, and $(\overline{Y},\overline{D})$ the toric surface with momentum polytope $\overline{P}$. The corresponding log Calabi–Yau surface $(Y,D)$ is obtained by blowing up four smooth points on the four components of the toric boundary divisor $\overline{D}$ of $\overline{Y}$. An associated Symington polygon $P$ with a polyhedral decomposition is illustrated in the middle of Figure 35 and corresponds to a one-parameter degeneration of the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}})$ to the union of two copies of $\mathbb{P}^{2}$. As described in §6.2, the mirror $\mathcal{X}$ to this one-parameter degeneration is a non-toric deformation of a toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$. It follows from the dual web represented on the right of Figure 35 that the central fiber $\mathcal{X}_{0}$ contains two compact irreducible components. The compact component corresponding to the quadrilatera in the web is isomorphic to $\mathbb{F}_{1}$. On the other hand, the compact component corresponding to the other bounded cell of the web is isomorphic to $\mathbb{F}_{6}$. The corresponding canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ is the total space of a smoothing of a cusp singularity $\overline{\mathcal{X}}_{0}^{\mathrm{can}}$. The associated rank two 5d SCFT appears as $\mathbb{F}_{1}\cup\mathbb{F}_{6}$ in [66, Figure 8], where it is also identified with the 5d $\mathcal{N}=1$ $\mathrm{Sp}(2)_{0}$ gauge theory.
Figure 35. The momentum polytope $\overline{P}$ for a polarized toric variety $\overline{Y}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}$ of $\overline{P}$ defining a maximal degeneration of $\overline{Y}$ in the middle, and the associated dual web of $5$-branes on the right, engineering a rank two 5d SCFT.
@@ 7.3. Relation to the Fanosearch program
By [33, Theorem 39], there exists actually a one-to-one correspondence between the set of $\mathbb{Q}$-Gorenstein degenerations of the Fano surface $X_{\overline{P}}$ to Fano toric surfaces and the set of Fano polygons obtained from $\overline{P}$ by mutations. Finally, one can attach to a Fano polygon $\overline{P}$ a log Calabi–Yau surface $(Y,D)$. For this, denote by $(\overline{Y},\overline{D},\overline{L})$ the polarized toric variety with momentum polytope $\overline{P}$. Then, $(Y,D)$ is obtained from $(\overline{Y},\overline{D})$ by blowing up for each elementary T-cone of $\overline{P}$ a point on the corresponding toric divisor of $\overline{Y}$. Finally, one can define a line bundle $L$ on $(Y,D)$ by $L=p^{\star}\overline{L}\otimes\mathcal{O}(-\sum_{i}a_{i}E_{i})$, where $p:Y\rightarrow\overline{Y}$ is the toric morphism, $E_{i}$ are the exceptional curves, and $a_{i}$ are the common lattice height/base of the corresponding elementary T-cones. The set of cluster charts $(\mathbb{C}^{\star})^{2}\subset U=Y\setminus D$ is in one-to-one correspondence with the set of mutations of $\overline{P}$, and the space of sections $H^{0}(Y,L)$ can be identified with the set of maximally mutable Laurent polynomials defined in [3, Definition 4], which are expected to be Landau-Ginzburg mirror to the Fano surface $X_{\overline{P}}$.
@@
The set-up of mirror symmetry for Fano surfaces briefly reviewed above can be viewed as a special case of the set-up considered in the present paper. Indeed, if $(Y,D,L)$ is a log Calabi–Yau surface with line bundle obtained from a Fano polygon $\overline{P}$, then, we have $L=\mathcal{O}(\sum_{i}a_{i}D_{i})$, where $D_{i}$ are the irreducible components of $D$, and $a_{i}$ is the lattice distance between the origin 0 and the side of $\overline{P}$ corresponding to $D_{i}$. In other words, we are in the particular case where the divisors defined by $L$ are linearly equivalent to a divisor supported on $D$. Conversely, if $(Y,D,L)$ is a log Calabi–Yau surface such that $L=\mathcal{O}(\sum_{i}a_{i}D_{i})$ with $a_{i}\in\mathbb{Z}_{\geq 0}$, then, for every toric model given by a polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ with momentum polytope $\overline{P}$, there exists a torus-invariant section of $\overline{L}$ with zero-divisor $\sum_{i}a_{i}\overline{D}_{i}$, which corresponds to a particular integral point in $\overline{P}$ which can be taken as the origin [77]. In this particular situation, the Fano polygon $\overline{P}$ naturally defines a Symington polygon $P$ by cutting out all elementary T-cones, so that the origin is the only singularity of the integral affine structure, see Figure 36. The corresponding non-compact Calabi–Yau 3-fold is the total space of the canonical line bundle $K_{X_{\overline{P}}}$ of the Fano surface $X_{\overline{P}}$. Indeed, starting from the toric Calabi–Yau 3-fold $K_{\overline{X}_{\overline{P}}}$, and applying the non-toric deformation described in §6.2, the Fano toric surface $\overline{X}_{\overline{P}}$ is deformed non-torically into the Fano surface $X_{\overline{P}}$. In particular, the corresponding canonical 3-fold singularity $\mathcal{X}^{\mathrm{can}}$ is obtained from $K_{X_{\overline{P}}}$ by contracting the zero section to a point. The 3-fold $K_{X_{\overline{P}}}$ is singular in general and so is only a partial crepant resolution of $\mathcal{X}^{\mathrm{can}}$. To obtain a crepant resolution, one should resolve the singularities by decomposing the Symington polygon $P$ into triangles of size one.
@@
Figure 36. On the left, a Fano polygon with a decomposition into elementary T-cones and R-cones. On the right, the corresponding Symington polygon obtained by cutting out the T-cones.
@@
As shown in [60], all $\mathbb{Q}$-Gorenstein degenerations of $\mathbb{P}^{2}$ are weighted projective spaces $\mathbb{P}(a^{2},b^{2},c^{2})$, for Markov triples $(a,b,c)$, that is, triple of positive integers such that $a^{2}+b^{2}+c^{2}=3abc$. Correspondingly, all mutation equivalent Fano polygons to the Fano polygon $\overline{P}$ in Example 7.1, can be characterized by a mutation sequence indexed by Markov triples as explained in [68, Example 2.4]. We illustrate in Figure 37 one of these mutation equivalent polygons $\overline{P}^{\prime}$, along with a polyhedral decomposition which determines a degeneration of the toric variety $(\overline{Y}^{\prime},\overline{D}^{\prime})$ with momentum polytope image $\overline{P}^{\prime}$. The mirror $\mathcal{X}\to\mathbb{C}$ to this degeneration is a deformation of the toric Calabi–Yau 3-fold $K_{\mathbb{P}(1,1,4)}$, and is still isomorphic to $K_{\mathbb{P}^{2}}$, as in Example 7.1. The compact component $\mathbb{P}^{2}$ of $\mathcal{X}_{0}$ is obtained from the compact component $\mathbb{P}(1,1,4)$ of $\overline{\mathcal{X}}_{0}$ by locally smoothing as in Equation (6.2) the elementary T-singularity $\mathbb{C}^{2}/(\mathbb{Z}/4\mathbb{Z})$ corresponding to the elementary T-cone drawn in red in Figure 37.
Figure 37. The momentum polytope $\overline{P}^{\prime}$ for a polarized toric variety $\overline{Y}^{\prime}$ on the left, a polyhedral decomposition $\overline{\mathcal{P}}^{\prime}$ of $\overline{P}^{\prime}$ defining a maximal degeneration of $\overline{Y}^{\prime}$ in the middle, and the associated dual web of $5$-branes on the right, engineering the rank one $E_{0}$ 5d SCFT.
Statements and declarations:
Conflict of interest statement: The authors have no competing interest, directly or indirectly related to this work, to declare.
Data availability statement: No data have been used or created for this work.